A company is considering an investment (at time = 0) in a wire and cable making machine. The cost of the machine is 44,741 dollars with zero expected salvage value. Annual production during the 3-year life of the machine is expected to be (starting at time = 1) 4,738, 8,973, and 13,807 units. The sale price per unit of wire is 13 dollars in year one, and then the price is expected to increase by 5% per year. Production costs per unit will be 5 dollars in year one, and then production costs are expected to increase by 4% per year. Depreciation on the machine is 12,217 dollars per year, the tax rate is 40% per year, and the minimum acceptable rate of return is 14% percent per year, compounded annually. Calculate the net present value of this investment. Assume all flows are at the end of each year. (note: round your answer to the nearest cent, and do not include spaces, currency signs, plus or minus signs, or commas)
Answer #1

Net cash flow = Annual revenue - Annual cost

Taxable income = Net cash flow - Depreciation

Tax = Tax rate * Taxable income

ATCF = Taxable income - Tax + Depreciation

Using Excel

Year Initial cost Revenue O&M cost Depreciation Salvage value TI Tax ATCF
0 -44741.00 -44741.00
1 61594.00 -23690.00 12217.00 25687.00 10274.80 22491.80
2 122481.45 -46659.60 12217.00 63604.85 25441.94 37658.94
3 197888.83 -74668.26 12217.00 0.00 111003.57 44401.43 56618.43
NPW 42181.81

Showing formula in excel

Year Initial cost Revenue O&M cost Depreciation Salvage value TI Tax ATCF
0 -44741 =B135
1 =4738*13 =-(5*4738) 12217 =C136+D136-E136+F136 =G136*0.4 =H136+E136
2 =13*1.05*8973 =-5*1.04*8973 12217 =C137+D137-E137+F137 =G137*0.4 =H137+E137
3 =13*1.05*1.05*13807 =-5*1.04*1.04*13807 12217 0 =C138+D138-E138+F138 =G138*0.4 =H138+E138
NPW =NPV(14%,I136:I138)+I135